# Scalable and Domain-General Abstractive Proposition Segmentation

### v1

@@ Abstract
Segmenting text into fine-grained units of meaning is important to a wide range of NLP applications. The default approach of segmenting text into sentences is often insufficient, especially since sentences are usually complex enough to include multiple units of meaning that merit separate treatment in the downstream task. We focus on the task of abstractive proposition segmentation: transforming text into simple, self-contained, well-formed sentences. Several recent works have demonstrated the utility of proposition segmentation with few-shot prompted LLMs for downstream tasks such as retrieval-augmented grounding and fact verification. However, this approach does not scale to large amounts of text and may not always extract all the facts from the input text. In this paper, we first introduce evaluation metrics for the task to measure several dimensions of quality. We then propose a scalable, yet accurate, proposition segmentation model. We model proposition segmentation as a supervised task by training LLMs on existing annotated datasets and show that training yields significantly improved results. We further show that by using the fine-tuned LLMs as teachers for annotating large amounts of multi-domain synthetic distillation data, we can train smaller student models with results similar to the teacher LLMs. We then demonstrate that our technique leads to effective domain generalization, by annotating data in two domains outside the original training data and evaluating on them. Finally, as a key contribution of the paper, we share an easy-to-use API for NLP practitioners to use.
@@ 1 Introduction
To provide useful fine-grained segmentation, several recent works have taken the approach of proposition segmentation [^1] Others in the literature have also used terms such as ”claim decomposition”, ”claim extraction”, and ”atomic fact extraction” for the same concept. We follow the naming in Chen et al. (2023b). Chen et al. (2023b); Min et al. (2023); Wanner et al. (2024), seeking to break text into fine-grained, minimal units of meaning that together convey all the information in the source text. Similarly to the extractive-vs-abstractive contrast in the summarization literature, the two strands of proposition segmentation work so far have considered either (a) an extractive approach, representing propositions as one or more spans in the source text (Chen et al. (2023b); Gunel et al. (2023); etc.); or (b) few-shot LLM prompts for abstractive proposition segmentation, generatively writing out each unit as a well-formed sentence Kamoi et al. (2023); Wanner et al. (2024).
@@
APS has already found applications in grounding Gao et al. (2023), summarization evaluation Liu et al. (2023), and fact checking Min et al. (2023). In this paper, we focus on making abstractive proposition segmentation practical. The few-shot prompting approaches are typically too costly to run at large scales, and, furthermore, we show that they tend to under-extract compared to our proposed solutions. Our core contributions are:
@@
    Scalable, domain-general student models (Gemma 7B, Mesnard et al. (2024)) for APS distilled from the supervised models over synthetic multi-domain data Hosseini et al. (2024), yielding performance comparable to the teacher models even on domains not seen in the human-annotated training data.
@@ 2 Related Work
Indeed, several previous results have shown APS by few-shot prompted LLMs benefits retrieval-augmented fact verification and grounding Kamoi et al. (2023); Min et al. (2023). A concurrent result in Wanner et al. (2024) looks more specifically at APS itself with few-shot prompting. They propose a specific single metric for APS, DecompScore, that combines our "reference-free precision" metric (§3.2) with the count of claims generated.
Other formats of proposition segmentation have also been explored. Extractive proposition segmentation is shown in Chen et al. (2023b, c) to benefit document-level NLI and retrieval. Several open-book QA and grounding works have generated fine-grained questions corresponding implicitly to the fine-grained claims in the text Gao et al. (2023); Chen et al. (2022, 2023a). In the summarization evaluation literature, "Summary Content Units", initially human-annotated Nenkova and Passonneau (2004), later generated from syntactic signals Gao et al. (2019) have long been used for summary evaluation. Before the LLM era, decomposing text into semantic triples, known as Open Information Extraction Etzioni et al. (2008), drove a variety of downstream applications.
Our desiderata for proposition segmentation include context-independence, earlier studied at the sentence level by Choi et al. (2021). It is tempting to analogize proposition segmentation to the work on text simplification Al-Thanyyan and Azmi (2021) , which makes text more accessible to humans. However, the desiderata of text simplification pertain to the limitations of human readers, and differ significantly from ours. In a sense, we can think of proposition segmentation as "text simplification for machine consumption" as contrasted against text simplification for human readers.
@@ 3.1 Task Definition
We are given an input text $t$, which comprises a naturally-occurring sequence of English words, possibly split into multiple sentences, i.e., $t=\{s_{1},...,s_{n}\}$. In text $t$, there are $k$ latent atomic facts $\{f_{1},...,f_{k}\}$. The task is then to segment $t$ into a list of propositions $\{p_{1},...,p_{k}\}$ with the following conditions:
@@
    Well-formed: Proposition $p_{i}$ should be grammatically correct and conform to the rules of the English language.
@@
    Atomic: Proposition $p_{i}$ should contain a single atomic fact.
@@
    Self-contained: Proposition $p_{i}$ should not need additional context to be understood.
@@
    Supported: Proposition $p_{i}$ should be found in the given text $t$.
@@
    Comprehensive: The list of propositions $\{p_{1},...,p_{k}\}$ should cover all the facts in text $t$.
@@ 3.2 Evaluation Metrics
To evaluate systems that produce propositions following the conditions above, we propose two sets of metrics that make use of an entailment model. We employ Natural Language Inference (NLI) as backbone to our metrics because by definition Dagan et al. (2013), it can be used to check factual support (i.e., one entails another) and semantic equivalence (i.e., both entail each other).
@@
where $\bar{Q}$ is the space-concatenated version of ${Q}$ to create a single text. Here, precision essentially evaluates whether each proposition $q_{i}$ in $Q$ is supported in text $t$, while recall evaluates whether each latent atomic fact mentioned in each sentence $s_{j}$ is covered in $Q$. We can then combine both precision and recall by calculating the f1-score $RF_{f1}$.
@@
The first NLI call (i.e., does gold entail predicted?) ensures atomicity: predicted proposition $q_{j}$ should not include other facts (otherwise, the entailment will not hold). The second NLI call (i.e., does predicted entail gold?) ensures self-containedness: $q_{j}$ should not need further context (otherwise, the entailment does not hold). We calculate precision $RB_{p}$ and recall $RB_{r}$ as follows:
@@ 4.2 Generating Multi-Domain Synthetic Data
We follow Hosseini et al. (2024) that take a practical approach and consider various text properties as contributing factors to domains: text genre, topic, and even the platform or venue that the text comes from. They design a prompt with $18$ few-shot examples, where each example is a triple of (length, domain, text). The length can take either the value short (just one or a few sentences) or paragraph. Appendix A shows an example. [^3] The full list can be found in Hosseini et al. (2024). The set of $18$ few-shot examples cover $8$ seed domains such as shopping reviews, twitter and reddit post. However, to have a wide range of domains, they first prompt FLAN-PaLM2 L (Unicorn) model Anil et al. (2023b) to generate new domains. Then, they manually select a number of non-repetitive domains. Finally, they prompt the LLM to generate text in those domains with the two lengths.
We replicated their approach using Gemini Ultra Anil et al. (2023a). We first prompted Gemini Ultra $4,000$ times to generate new domains. We obtained a set of $105$ domains, from which we manually selected $75$. We then prompted the LLM and generated $226$K examples with the selected domains and the two lengths. [^4] We first generated $228$K examples, but filtered examples with $n\geq 4$-gram overlap with any of the seed examples ($\approx 2$K examples).
@@ 4.3 Distillation
In our preliminary experiments, we observed better results from the grouped propositions version (§5.3), so we trained the student model based on this type of teacher. We apply the teacher LLMs to the synthetic multi-domain set of texts (§4.2) and produce $226$K (text, propositions) pairs. We then train a model with the same input and output format as the teacher models with grouped proposition. We used Gemma 7b Mesnard et al. (2024), a lightweight state-of-the-art LM, as our student model.
@@ 5 Experiments
5.1 Datasets
ROSE dataset. This dataset is built by manually splitting news summaries into Atomic Content Units (ACUs) for the purpose of evaluating such summaries Liu et al. (2023). The ACUs in this dataset are very close to our propositions definition, therefore we used them for training. We observed some cases in the dataset where the propositions are either not supported or not comprehensive, but we filtered those examples automatically. The dataset contains $2,500$ passages ($21,797$ propositions). We randomly split the dataset into a training and development set (for hyper-parameter tuning [^5] See details of hyper-parameters in Appendix B.). The training set contains $2,089$ passages ($18,994$ propositions), and the development set contains $411$ passages ($2,803$ propositions).
The original dataset contains the full set of propositions for each passage. However, for training the grouped propositions version (§4.1), we need to align each proposition to a sentence. We pre-process the dataset to obtain such alignment. We use the NLI score (from T5 11B trained on ANLI) between sentences (premise) and propositions (hypothesis) to obtain the alignment. In particular, for each proposition, we find the sentence with the maximum NLI score. If the NLI score from that sentence $\geq\tau=0.9$, we align the proposition to the sentence. Otherwise, we discard the example (unsupported proposition). After aligning all the propositions, if a sentence is not aligned with any proposition, we again discard the example (a special case of non-comprehensive propositions). We provide more details about the alignment and filtering in Appendix C.
5.2 Baselines
@@
Few Shot extracts propositions by few-shot prompting an LLM. For each test example, we selected the most similar $K=10$ examples from the training set based on ROUGE-1 score Lin (2004). We report the results for two LLMs, Gemini Pro and Gemini Ultra. We also tried two additional few-shot prompting approaches: A) Randomly sampling few-shot examples (average of $5$ runs), B) The few-shot examples from Wanner et al. (2024). In both cases, the results were overall worse than the dynamic approach and had a similar pattern compared to our models (Appendeix D).
Trained on ROSE are cases where we supervised trained a LM. We trained two versions for each language model (§4.1): ungrouped propositions (UG) and grouped propositions (G). We trained Gemma 7B, Gemini Pro and Gemini Ultra. We also tried Gemma 2b and obtained consistent results with 7b (Appendix D). Moreover, we did preliminary experiments with T5 and obtained consistent results, although lower than Gemma.
@@ 5.3 In-Domain Results
Student models. We trained two different Gemma 7B student models, one trained on distillation data from Gemini Pro teacher model, and one from Gemini Ultra teacher model. The Gemma 7B student models outperform Gemma 7B trained directly on ROSE on all metrics. In addition, Gemma 7B student models (the last two rows) perform close to their corresponding teacher models (the last two rows of the trained on ROSE block).
@@ 5.4 Out-of-Domain Results
Student models vs teacher models and training directly on ROSE. In both datasets, all the trained and distilled models have very high $RL$ metrics ($RL_{f1}\geq 96\%$). However, the student models perform significantly better than Gemma 7B trained directly on ROSE in $RB$ metrics. This confirms that our distillation approach using synthetic multi-domain data leads to successful domain adaptation. In addition, the student models get results on par with teacher models (and sometimes even better) on out-of-domain datasets.
Few-shot models compared to student and teacher. The few-shot models have very low $RL_{r}$ ($53\%$ to $83\%$) compared to the student models ($\geq 97$%). This makes the few-shot models unreliable for downstream applications such as fact verification that require to have access to all the claims in the input passage. Table 11 in Appendix E shows examples. The $RB$ metrics for few-shot models is slightly better than the student models on Reddit, but much worse on Amazon Review.
@@
In many cases, it is challenging and subjective to decide on the right level of atomicity and decontextualization, which directly affects $RB$ metrics (§3.2). Table 4 shows a number of examples where our annotated and predicted propositions (Gemma 7b distilled from Gemini Pro data) are different, although the predicted ones are not necessarily wrong especially when it comes to the atomicity level. For example, the sentence “There are lots of things that feel good that carry some kind of risk” has the right level of atomicity to be considered as a proposition if we want to emphasize on the two points jointly (“feeling good” and “carrying some kind of risk”). Otherwise, the sentence could be split into two propositions.

### v2

@@ Abstract
Segmenting text into fine-grained units of meaning is important to a wide range of NLP applications. The default approach of segmenting text into sentences is often insufficient, especially since sentences are usually complex enough to include multiple units of meaning that merit separate treatment in the downstream task. We focus on the task of abstractive proposition segmentation (APS): transforming text into simple, self-contained, well-formed sentences. Several recent works have demonstrated the utility of proposition segmentation with few-shot prompted LLMs for downstream tasks such as retrieval-augmented grounding and fact verification. However, this approach does not scale to large amounts of text and may not always extract all the facts from the input text. In this paper, we first introduce evaluation metrics for the task to measure several dimensions of quality. We then propose a scalable, yet accurate, proposition segmentation model. We model proposition segmentation as a supervised task by training LLMs on existing annotated datasets and show that training yields significantly improved results. We further show that by using the fine-tuned LLMs (Gemini Pro and Gemini Ultra) as teachers for annotating large amounts of multi-domain synthetic distillation data, we can train smaller student models (Gemma 1 2B and 7B) with results similar to the teacher LLMs. We then demonstrate that our technique leads to effective domain generalization, by annotating data in two domains outside the original training data and evaluating on them. Finally, as a key contribution of the paper, we share an easy-to-use API [^1] Our Gemma-APS API (Gemma 1 2B and 7B) can be found on Hugging Face. for NLP practitioners to use.
@@ 1 Introduction
To provide useful fine-grained segmentation, several recent works have taken the approach of proposition segmentation [^2] Others in the literature have also used terms such as “claim decomposition”, “claim extraction”, and “atomic fact extraction” for the same concept. We follow the naming in Chen et al. (2023b). Chen et al. (2023b); Min et al. (2023); Wanner et al. (2024), seeking to break text into fine-grained, minimal units of meaning that together convey all the information in the source text. Similarly to the extractive-vs-abstractive contrast in the summarization literature, the two strands of proposition segmentation work so far have considered either (a) an extractive approach, representing propositions as one or more spans in the source text (Chen et al. (2023b); Gunel et al. (2023); etc.); or (b) few-shot LLM prompts for abstractive proposition segmentation, generatively writing out each unit as a well-formed sentence Kamoi et al. (2023); Wanner et al. (2024); Scirè et al. (2024).
@@
APS has already found applications in grounding Gao et al. (2023a), summarization evaluation Liu et al. (2023); Scirè et al. (2024), and fact checking Min et al. (2023). In this paper, we focus on making abstractive proposition segmentation practical. The few-shot prompting approaches are typically too costly to run at large scales, and, furthermore, we show that they tend to under-extract compared to our proposed solutions. Our core contributions are:
@@
    Scalable, domain-general student models (Gemma 1 2B and 7B, Mesnard et al. (2024)) for APS distilled from the supervised models over synthetic multi-domain data Hosseini et al. (2024), yielding performance comparable to the teacher models even on domains not seen in the human-annotated training data.
@@ 2 Related Work
Indeed, several previous results have shown APS by few-shot prompted LLMs benefits retrieval-augmented fact verification and grounding Kamoi et al. (2023); Min et al. (2023). A concurrent result in Wanner et al. (2024) looks more specifically at APS itself with few-shot prompting. Scirè et al. (2024) also perform few-shot prompting followed by distillation.
Other formats of proposition segmentation have also been explored. Extractive proposition segmentation is shown in Chen et al. (2023b, c) to benefit document-level NLI and retrieval. Several open-book QA and grounding works have generated fine-grained questions corresponding implicitly to the fine-grained claims in the text Gao et al. (2023a); Chen et al. (2022, 2023a). In the summarization evaluation literature, “Summary Content Units”, initially human-annotated Nenkova and Passonneau (2004), later generated from syntactic signals Gao et al. (2019) have long been used for summary evaluation. Before the LLM era, decomposing text into semantic triples, known as Open Information Extraction Etzioni et al. (2008), drove a variety of downstream applications.
Our desiderata for proposition segmentation include context-independence, earlier studied at the sentence level by Choi et al. (2021). Deng et al. (2024) perform document-level claim extraction for fact checking. They specifically extract claims that are check-worthy, where these claims are decontextualized, but not necessarily atomic.
In our work, we propose a suite of automatic evaluation metrics. Previous efforts have not focused much on metric definition. The exceptions are two concurrent works: A) Wanner et al. (2024) propose a specific single metric for APS, DecompScore, that combines our reference-free precision metric (§3.2) with the count of claims generated. B) Scirè et al. (2024) propose metrics based on ROUGE Lin (2004) and similar to our reference-based precision and recall. Our metrics are based on NLI that is more suitable for checking semantic equivalence of predicted and gold propositions.
@@ 3.1 Task Definition
We are given an input text $t$, which comprises a naturally-occurring sequence of English words, possibly split into multiple sentences, i.e., $t=\{s_{1},...,s_{n}\}$. In text $t$, there are $k$ latent gold propositions (claims) $\{p_{1},...,p_{k}\}$. The task is then to segment $t$ into a list of propositions $\{q_{1},...,q_{k}\}$ with the following conditions:
@@
    Well-formed: Proposition $q_{i}$ should be grammatically correct and conform to the rules of the English language.
@@
    Atomic: Proposition $q_{i}$ should contain a single atomic fact.
@@
    Self-contained: Proposition $q_{i}$ should not need additional context to be understood.
@@
    Supported: Proposition $q_{i}$ should be found in the given text $t$.
@@
    Comprehensive: The list of propositions $\{q_{1},...,q_{k}\}$ should cover all the latent gold propositions (claims) in text $t$.
@@ 3.2 Evaluation Metrics
To evaluate systems that produce propositions following the conditions above, we propose two sets of metrics that make use of an entailment model. We employ Natural Language Inference (NLI) as backbone to our metrics because by definition Dagan et al. (2013), it can be used to check factual support (i.e., one entails another) and semantic equivalence (i.e., both entail each other). In addition, NLI has been successfully used for evaluating factual consistency Anil et al. (2023a); Gao et al. (2023b); Fierro et al. (2024); Honovich et al. (2022).
@@
where $\bar{Q}$ is the space-concatenated version of ${Q}$ to create a single text. Here, precision essentially evaluates whether each proposition $q_{i}$ in $Q$ is supported in text $t$, while recall evaluates whether each latent gold proposition mentioned in each sentence $s_{j}$ is covered in $Q$. We can then combine both precision and recall by calculating the f1-score $RF_{f1}$.
@@
The first NLI call (i.e., does gold entail predicted?) ensures atomicity: If the predicted proposition $q_{j}$ is not as atomic as a gold proposition $p_{i}$, then $p_{i}$ will not entail $q_{j}$ (since $q_{j}$ has more information that $p_{i}$). The second NLI call (i.e., does predicted entail gold?) ensures self-containedness: If the predicted proposition $q_{j}$ is not as self-contained as a gold proposition $p_{i}$, then $q_{j}$ will not entail $p_{i}$ (since $p_{i}$ has more information than $q_{j}$).
$q_{j}$ should not need further context (otherwise, the entailment does not hold). We calculate precision $RB_{p}$ and recall $RB_{r}$ as follows:
@@
In order to validate the effectiveness of our metrics, we perform human correlation studies and show positive results (§6). We show examples of how metrics are calculated in Appendix A.
@@ 4.2 Generating Multi-Domain Synthetic Data
We follow Hosseini et al. (2024) that take a practical approach and consider various text properties as contributing factors to domains: text genre, topic, and even the platform or venue that the text comes from. They design a prompt with $18$ few-shot examples, where each example is a triple of (length, domain, text). The length can take either the value short (just one or a few sentences) or paragraph. Appendix B shows an example. [^4] The full list can be found in Hosseini et al. (2024). The set of $18$ few-shot examples cover $8$ seed domains such as shopping reviews, twitter and reddit post. However, to have a wide range of domains, they first prompt FLAN-PaLM2 L (Unicorn) model Anil et al. (2023b) to generate new domains. Then, they manually select a number of non-repetitive domains. Finally, they prompt the LLM to generate text in those domains with the two lengths.
We replicated their approach using Gemini Ultra Anil et al. (2023a). We first prompted Gemini Ultra $4,000$ times to generate new domains. [^5] Many of the calls generate one of the existing domains from the few-shot examples. Therefore, in order to obtain many unseen domains, we prompted the LLM $4,000$ times. We obtained a set of $105$ domains, from which we manually selected $75$. We then prompted the LLM and generated $226$K examples with the selected domains and the two lengths. [^6] We first generated $228$K examples, but filtered examples with $n\geq 4$-gram overlap with any of the seed examples ($\approx 2$K examples).
@@ 4.3 Distillation
In our preliminary experiments, we observed better results from the grouped propositions version (§5.3), so we trained the student model based on this type of teacher. We apply the teacher LLMs to the synthetic multi-domain set of texts (§4.2) and produce $226$K (text, propositions) pairs. We then train a model with the same input and output format as the teacher models with grouped proposition. We used Gemma 1777We used Gemma 1 in all our experiments, but we refer to the language model as Gemma for simplicity throughout most of the following text. (2B and 7B) Mesnard et al. (2024), a lightweight state-of-the-art LM [^8] https://ai.google.dev/gemma, as our student model.
@@ 5 Experiments
5.1 Datasets
ROSE dataset. This dataset is built by manually splitting news summaries into Atomic Content Units (ACUs) for the purpose of evaluating such summaries Liu et al. (2023). The ACUs are annotated based on a set of well-defined rules to extract atomic facts, i.e., elementary information units in the input text which no longer need to be further split for the purpose of ambiguity reduction for human evaluation Liu et al. (2023).
The ACU definition in the ROSE dataset are very close to our propositions definition, therefore we used them for training. We observed some cases in the dataset where the propositions are either not supported or not comprehensive, but we filtered those examples automatically. The dataset contains $2,500$ passages ($21,797$ propositions). We randomly split the dataset into a training and development set (for hyper-parameter tuning [^9] See details of hyper-parameters in Appendix C.). The training set contains $2,089$ passages ($18,994$ propositions), and the development set contains $411$ passages ($2,803$ propositions).
The original dataset contains the full set of propositions for each passage. However, for training the grouped propositions version (§4.1), we need to align each proposition to a sentence. We pre-process the dataset to obtain such alignment. We use the NLI score (from T5 11B trained on ANLI) between sentences (premise) and propositions (hypothesis) to obtain the alignment. In particular, for each proposition, we find the sentence with the maximum NLI score. If the NLI score from that sentence $\geq\tau=0.9$, we align the proposition to the sentence. Otherwise, we discard the example (unsupported proposition). After aligning all the propositions, if a sentence is not aligned with any proposition, we again discard the example (a special case of non-comprehensive propositions). We provide more details about the alignment, filtering, and pre-processing in Appendix D.
Since the examples in the ROSE dataset are based on news summaries, they are quite general and cover many linguistic forms such as presuppositions, attribution to the speaker, modals, and sentence connectors (e.g., because and however).
The manual annotations for the Reddit and Amazon Review datasets were done by two of the authors (each annotated one dataset). The instructions were based on the task definition (§3.1). The authors looked at examples from the ROSE dataset to be mostly aligned with those examples as well.
5.2 Baselines
@@
Few Shot extracts propositions by few-shot prompting an LLM. For each test example, we selected the most similar $K=10$ examples from the training set based on ROUGE-1 score Lin (2004). We report the results for two LLMs, Gemini Pro and Gemini Ultra. We also tried two additional few-shot prompting approaches: A) Randomly sampling few-shot examples (average of $5$ runs), B) The few-shot examples from Wanner et al. (2024). In both cases, the results were overall worse than the dynamic approach and had a similar pattern compared to our models (Appendeix E).
Trained on ROSE are cases where we supervised trained a LM. We trained two versions for each language model (§4.1): ungrouped propositions (UG) and grouped propositions (G). We trained Gemma 7B, Gemini Pro and Gemini Ultra. We also tried Gemma 2B and obtained consistent results with 7B (Appendix E). Moreover, we did preliminary experiments with T5 and obtained consistent results, although lower than Gemma.
@@ 5.3 In-Domain Results
Student models. We trained two different Gemma 7B student models, one trained on distillation data from Gemini Pro teacher model, and one from Gemini Ultra teacher model. The Gemma 7B student models outperform Gemma 7B trained directly on ROSE (i.e., with no distillation) on all metrics. In addition, Gemma 7B student models (the last two rows) perform close to their corresponding teacher models (the last two rows of the trained on ROSE block).
@@ 5.4 Out-of-Domain Results
Student models vs teacher models and training directly on ROSE. In both datasets, all the trained and distilled models have very high $RL$ metrics ($RL_{f1}\geq 96\%$). However, the student models perform significantly better than Gemma 7B trained directly on ROSE (i.e., with no distillation) in $RB$ metrics. This confirms that our distillation approach using synthetic multi-domain data leads to successful domain adaptation. In addition, the student models get results on par with teacher models (and sometimes even better) on out-of-domain datasets.
@@
Few-shot models compared to student and teacher. The few-shot models have very low $RL_{r}$ ($53\%$ to $83\%$) compared to the student models ($\geq 97$%). This makes the few-shot models unreliable for downstream applications such as fact verification that require to have access to all the claims in the input passage. Table 13 in Appendix F shows examples. The $RB$ metrics for few-shot models is slightly better than the student models on Reddit, but much worse on Amazon Review.
@@
In many cases, it is challenging and subjective to decide on the right level of atomicity and decontextualization, which directly affects $RB$ metrics (§3.2). Table 4 shows a number of examples where our annotated and predicted propositions (Gemma 7B distilled from Gemini Pro data) are different, although the predicted ones are not necessarily wrong especially when it comes to the atomicity level. For example, the sentence “There are lots of things that feel good that carry some kind of risk” has the right level of atomicity to be considered as a proposition if we want to emphasize on the two points jointly (“feeling good” and “carrying some kind of risk”). Otherwise, the sentence could be split into two propositions. In our work, since we trained the teacher LLMs on the ROSE dataset, we expect the trained models to mirror the atomicity levels in the dataset.
6 Human Correlation Studies for APS Metrics
We measure how well our defined metrics (§3.2) align with human judgements in order to validate the metrics’ effectiveness. We performed a study on $40$ passages ($142$ sentences, $263$ predicted propositions, and $262$ gold propositions) from the Amazon Review dataset. We used the predictions from two models: Gemini Pro few-shot and Gemma 7B distilled from fine-tuned Gemini Pro. The annotations were done by two of the authors (each example were annotated by one author). For each input sentence, we annotated whether the predicted propositions cover all the claims in the sentence (used for measuring reference-less recall). For each predicted proposition, we annotated whether it is supported by the input passage (reference-less precision) and whether it is equivalent to any of the gold propositions (reference-based precision). For each gold proposition, we annotated whether it is equivalent to any of the predicted propositions (reference-based recall).
The Pearson correlation coefficients of example-level metrics and human annotations were generally high (Table 5), confirming that our proposed metrics do correlate well with human judgements (p-value < $0.01$). [^11] The reference-less precision metric is almost always equal to 1 except for a few examples. The NLI accuracy compared to human annotation is $0.985$. In addition, both the automatic metric and the human evaluated metric are $>=0.98$.
Metric Pearson Correlation Coefficient
Reference-based Pr 0.718
Reference-based Rec 0.731
Reference-less Pr 0.476
Reference-less Rec 0.647
Table 5: Pearson correlation coefficients of metrics and human judgements (p-value < $0.01$).